Two charges equal in magnitude and opposite in polarity are placed at certain distance apart and force acting between them is F. If 75 % charge of one is transferred to another, then the force between charges becomes

Grade:12

1 Answers

Khimraj
6 years ago
Just write the expression and then compareF= Kq(-q)/r²Now q` = 1/4qAnd -q`= -1/4qF` = k(q/4)(-q/4)/r²F` = F/16